OpenStudy (anonymous):

To skate at Roller Heaven, each person must pay a membership fee (m) and a fee for each session (f). Art attended 5 sessions and paid $25 in all. Later, when he had attended 10 sessions, he paid $40 in all. How much is the membership fee and how much are the sessions. Show how to solve this, using any method.

OpenStudy (anonymous):

posting the same question twice strange......

OpenStudy (anonymous):

one method is finding the equation for the line between (5, 25) and (10,40)

OpenStudy (anonymous):

it's not, hashir! it's sifferent!

OpenStudy (anonymous):

or we can just reason it out

OpenStudy (anonymous):

*different

OpenStudy (anonymous):

you pick

OpenStudy (anonymous):

so m+5f=25 and m+10f=40

OpenStudy (sriram):

m+5f=25 m+10f=40 f=3 m=10

OpenStudy (anonymous):

its the same one!!!!!!!!!!

OpenStudy (anonymous):

yikes. ok we reason it out. 5 sessions cost $25 10 cost $40 so the 5 additional sessions cost an additional $15 therefore each one cost $3

OpenStudy (anonymous):

since you know each one cost $3 then 5 sessions cost $15 for the sessions and therefore since the total was $25 the flat fee must have been $10

OpenStudy (anonymous):

this is exactly the same as saying "find the equation for the line between the points (5,25) and (10,40) the "slope" of the line is \[\frac{40-25}{10-5}=\frac{15}{5}=3\]

OpenStudy (anonymous):

that is the rate per class. the y - intercept is 10, that is the flat fee. you get it by solving \[25=3\times 5+b\] for b